The following simple random sample was selected from a normal distribution: 4, 6, 3, 5, 9, 3.



1. Construct a 90% confidence interval for the population mean μ.




2. Construct a 95% confidence interval for the population mean μ.




3. Construct a 99% confidence interval for the population mean μ.




4. Assume that the sample mean x and sample standard deviation s remain exactly the same as those you just calculated but that they are based on a sample of n = 25 observations rather than n = 6 observations. Repeat parts 1-3. What is the effect of increasing the sample size on the width of the confidence intervals?




1
Expert's answer
2021-11-15T01:44:34-0500
"mean=\bar{x}=\dfrac{1}{6}(4+6+3+5+9+3)=5"

"s^2=\dfrac{1}{6-1}((4-5)^2+(6-5)^2+(3-5)^2+(5-5)^2"

"+(9-5)^2+(3-5)^2)=5.2"




"s=\sqrt{s^2}=\sqrt{5.2}\approx2.280351"

1. The critical value for "\alpha = 0.1" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.015036." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-2.015036\times\dfrac{2.280351}{\sqrt{6}},"

"5+2.015036\times\dfrac{2.280351}{\sqrt{6}})"

"=(3.1241, 6.8759)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "3.1241 < \mu < 6.8759," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(3.1241, 6.8759)."


2. The critical value for "\alpha = 0.05" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.570543." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-2.570543\times\dfrac{2.280351}{\sqrt{6}},"

"5+2.570543\times\dfrac{2.280351}{\sqrt{6}})"

"=(2.6070, 7.3930)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "2.6070 < \mu < 7.3930," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(2.6070, 7.3930)."


3. The critical value for "\alpha = 0.01" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =4.031677." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-4.031677\times\dfrac{2.280351}{\sqrt{6}},"

"5+4.031677\times\dfrac{2.280351}{\sqrt{6}})"

"=(1.2467, 8.7533)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "1.2467 < \mu < 8.7533," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(1.2467, 8.7533)."


4.

a. The critical value for "\alpha = 0.1" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =1.710882." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-1.710882\times\dfrac{2.280351}{\sqrt{25}},"

"5+1.710882\times\dfrac{2.280351}{\sqrt{25}})"

"=(4.2197, 5.7803)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "4.2197 < \mu < 5.7803," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(4.2197, 5.7803)."


b. The critical value for "\alpha = 0.05" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.063899." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-2.063899\times\dfrac{2.280351}{\sqrt{25}},"

"5+2.063899\times\dfrac{2.280351}{\sqrt{25}})"

"=(4.0587, 5.9413)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "4.0587 < \mu < 5.9413," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(4.0587, 5.9413)."


c. The critical value for "\alpha = 0.01" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.79694." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5-2.79694\times\dfrac{2.280351}{\sqrt{25}},"

"5+2.79694\times\dfrac{2.280351}{\sqrt{25}})"

"=(3.7244, 6.2756)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "3.7244 < \mu < 6.2756," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(3.7244, 6.2756)."


The width of a confidence interval decreases as the sample size increases and increases as the confidence level increases.

Larger samples give narrower intervals. We are able to estimate a population mean more precisely with a larger sample size.